# UnLoc: Leveraging Depth Uncertainties for Floorplan Localization

### v1

@@ Abstract
We propose UnLoc, an efficient data-driven solution for sequential camera localization within floorplans. Floorplan data is readily available, long-term persistent, and robust to changes in visual appearance. We address key limitations of recent methods, such as the lack of uncertainty modeling in depth predictions and the necessity for custom depth networks trained for each environment. We introduce a novel probabilistic model that incorporates uncertainty estimation, modeling depth predictions as explicit probability distributions. By leveraging off-the-shelf pre-trained monocular depth models, we eliminate the need to rely on per-environment-trained depth networks, enhancing generalization to unseen spaces. We evaluate UnLoc on large-scale synthetic and real-world datasets, demonstrating significant improvements over existing methods in terms of accuracy and robustness. Notably, we achieve $2.7$ times higher localization recall on long sequences (100 frames) and $16.7$ times higher on short ones (15 frames) than the state of the art on the challenging LaMAR HGE dataset.
Figure 1: The proposed UnLoc processes an input image sequence to predict the floorplan depth (in meters) and associated uncertainty for each image column. Using these predictions, it generates a probability distribution over potential $\text{SE}(2)$ camera poses and outputs the most likely one (blue arrow). The ground truth pose of the last frame is also shown (red arrow), overlapped by the predicted pose.
@@ 1 Introduction
Camera localization within indoor environments is a fundamental problem in computer vision, essential for applications in augmented reality and robotics. Accurate localization enables devices to understand their spatial context, facilitating tasks such as navigation, object interaction, and autonomous exploration. Traditional localization methods often rely on pre-built 3D models [41, 43, 30, 39, 36] or extensive image databases [44, 2, 23, 48], which are storage-intensive and require substantial maintenance, limiting their scalability to new or dynamically changing environments.
@@
Floorplans offer a lightweight and readily available alternative for indoor localization. As 2D representations of spaces, floorplans are easy to obtain [29, 56] and remain unaffected by changes in appearance, such as furniture rearrangements or lighting variations. Recent methods have leveraged floorplans for localization by aligning images with the map [18, 32, 10], enabling devices to localize within new environments as long as a floorplan is available.
@@
Among these, F3Loc [10] has emerged as a recent promising approach for sequential visual floorplan localization, significantly outperforming all previous baselines. It integrates observations over time using a histogram filter, achieving impressive accuracy and computational efficiency on a range of datasets. However, as promising as it is, F3Loc has notable limitations that hinder its practical deployment, which we aim to improve upon in this paper:
@@
Lack of Uncertainty Modeling: F3Loc combines monocular and multi-view depth estimation. However, it assumes that the resulting predictions are all of similar accuracy, and it has no means to represent and predict the uncertainties in the depth predictions. In indoor environments, depth estimation is often unreliable in regions with glass walls, open doorways, or large, featureless walls. When fusing a sequence of predictions, not accounting for uncertainty, inaccurate depth predictions adversely affect the localization process, leading to erroneous pose estimates.
@@
Dataset-Specific Depth Networks: F3Loc relies on a custom depth prediction network that must be trained separately for each dataset or environment. This per-dataset training requirement poses significant challenges for scalability and robustness. Collecting sufficient depth data for retraining the depth network for every new environment is impractical, especially when rapid deployment is desired.
@@
This paper addresses these shortcomings by introducing a novel visual floorplan localization method that incorporates uncertainty estimation into monocular depth prediction and uses it to robustly fuse predictions from a sequence of images. Also, our method enables us to leverage pre-trained monocular depth models. Our contributions are:
@@
  • Uncertainty-Aware Depth Prediction: We model floorplan depth predictions as explicit probability distributions, enabling us to represent the uncertainty associated with each estimate. This uncertainty-aware approach improves localization by weighting predictions according to their confidence levels, which is particularly beneficial in challenging areas prone to depth estimation errors.
  • Leveraging Off-the-Shelf Depth Models: Instead of training custom depth networks for each environment, we utilize state-of-the-art monocular depth networks pre-trained on large-scale datasets [53]. These models provide accurate depth predictions out-of-the-box, eliminating the need for extensive per-environment training and enhancing the method’s generalization to new settings.
@@
The proposed UnLoc (see Fig. 1) demonstrates significant improvements in accuracy and robustness on multiple datasets compared to the state of the art.
@@ 2 Related work
Figure 2: Method overview. At timestep $t$, UnLoc aligns an image with gravity and processes it through a pretrained monodepth encoder. Features extracted at this step, along with a binary mask from the gravity alignment, are used to predict the floorplan depth $\mathbf{\hat{d}}_{t}$ and its uncertainty $\mathbf{\hat{b}}_{t}$ via a masked attention mechanism. These predictions are used to form equiangular rays to perform uncertainty-aware matching with the floorplan’s occupancy map. Finally, a histogram filter fuses the observation likelihood with the integrated past belief.
@@
Visual localization is a fundamental problem in computer vision, addressed through various approaches. Traditional methods include image retrieval techniques [12, 21, 2, 23, 48], which find the most similar images in a database and estimate the pose of the query image based on the retrieved ones. Structure-from-Motion-based approaches [1, 44, 42, 36] build a 3D model of the environment and establish 2D-3D correspondences by matching local descriptors, computing camera poses using minimal solvers [27] and RANSAC [15] or its recent variants [4, 3]. Scene coordinate regression methods [8, 9] learn to regress the 3D coordinates of image pixels, while pose regression techniques [25, 24] use networks to predict a 6-DoF camera pose from input images directly. These methods often rely on pre-built 3D models that are storage-intensive and scene-specific, limiting their applicability in unseen environments.
@@
To overcome this, floorplan-based localization methods have emerged, utilizing overhead images or floorplans to estimate the $\text{SE}(2)$ pose of the camera [50, 46]. These approaches can localize images in new scenes as long as a floorplan is provided. Floorplan localization is frequently associated with LiDAR sensors [33, 54, 58], which are impractical for widespread mobile device use. Alternative methods reconstruct 3D geometry using depth cameras [49] or visual odometry [34]. Some approaches extract geometric features like room edges to align with the floorplan [7, 28]. However, these methods often assume known camera or room height, which is not always feasible.
@@
Recent learning-based methods aim to use only RGB images for floorplan localization. LaLaLoc [19] estimates the position of panoramic images in a floorplan by embedding map and image features into a shared space. LaLaLoc++ [18] removes the known camera and ceiling height assumption by directly embedding the floorplan. LASER [32] represents the floorplan as a set of points and uses PointNet [37] to embed the visible points for each pose, aligning them with image features in a shared space. PF-Net [22] integrates localization within a differentiable particle filtering framework, using a learned similarity between images and corresponding map patches. F3Loc [10] utilizes metric monocular depth prediction but requires training custom depth networks, which can limit generalizability and assumes that all predictions are of the same quality.
@@
Sequence-based localization methods enhance robustness by integrating information over time. Bayesian filtering [13, 11, 22, 7, 31] is commonly used to fuse sequential observations, like particle [13] and histogram filters [45]. PF-Net [22] employs a differentiable particle filter for localization but relies on learned observation models that may not generalize well. F3Loc [10] uses sequential observations and integrates them employing a histogram filter.
@@
Depth estimation provides valuable geometric information for localization [10]. Nowadays, there are many strong pre-trained monodepth models [38, 5, 16, 55, 20, 52, 53, 6], trained on large-scale datasets, which can provide accurate metric and/or relative depth predictions out-of-the-box without the need for per-scene training. In this work, we leverage off-the-shelf monodepth networks with uncertainty estimation to enhance image-based floorplan localization. By utilizing pre-trained monodepth networks, we avoid custom depth model training, improving generalization. Incorporating uncertainty estimation enables us to guide the localization process using the confidence of depth predictions, focusing on reliable areas and improving accuracy.
@@ 3 Floorplan Localization with UnLoc
3.1. Method Overview. We estimate the 2D pose $\mathbf{s}_{t}=[s_{x,t},s_{y,t},s_{\phi,t}]$ within a known 2D floorplan, where $s_{x,t}$ and $s_{y,t}$ denote position coordinates and $s_{\phi,t}$ represents orientation. Our approach takes as input a sequence of $t$ RGB images, relative poses between these images, gravity direction, camera intrinsics, and the geometric layout of the floorplan. To ensure out-of-the-box usability of the method, the input floorplans are provided solely as occupancy grid data without any semantic labels.
An overview of our method is in Fig. 2. Each image is aligned with gravity and processed by a pre-trained encoder [53] to extract features. These features, along with a binary mask from the gravity alignment, are input to a masked attention mechanism that predicts the floorplan depth, defined as the depth to the nearest occupied area in the floorplan. The predicted depth is then used to construct equiangular rays, which are matched against the floorplan to produce a localization estimate. Finally, we combine this estimate with the previous posterior localization using a histogram filter to compute the updated posterior estimate.
@@
3.2. Gravity Alignment. Real-world applications often involve images captured from hand-held or head-mounted devices, resulting in arbitrary camera orientations. To address this, we preprocess images to align them with the gravity.
@@
We achieve gravity alignment by utilizing the camera’s roll ($\psi$) and pitch ($\theta$) angles to define rotation matrices. The rotation matrix for roll is given as $\mathbf{R}_{x}(\psi)$ and for pitch as $\mathbf{R}_{y}(\theta)$. The combined rotation matrix from the gravity-aligned frame to the camera frame is calculated as $\mathbf{R}_{cg}=\mathbf{R}_{y}(\theta)\cdot\mathbf{R}_{x}(\psi)$. The inverse transformation, from the camera frame to the gravity-aligned frame, is given by: $\mathbf{R}_{gc}=\mathbf{R}_{cg}^{\top}$. Using these rotations, we compute a homography $\mathbf{H}$ to warp the image into the gravity-aligned frame $\mathbf{H}=\mathbf{K}\cdot\mathbf{R}_{gc}\cdot\mathbf{K}^{-1}$, where $\mathbf{K}$ is the camera intrinsic matrix.
This process also produces a binary mask indicating pixels that are invalid after alignment due to the warping. The gravity-aligned RGB image and the mask are then used for subsequent feature extraction.
@@
Note that the gravity direction and camera intrinsics are usually accessible from sensors on smartphones and head-mounted devices. If this information is not directly available, methods such as GeoCalib [47] can be used to estimate both the camera intrinsics and gravity direction.
@@
3.3. Feature Extraction. Recent visual floorplan localization methods [19, 18, 32, 10] rely on encoders pre-trained on ImageNet [14] for image classification tasks, such as ResNet-50 [17]. However, we posit that encoders trained on tasks more closely related to floorplan depth estimation could yield better performance. To explore this, we utilize encoders pre-trained on dense monocular depth estimation tasks. These encoders, optimized on large-scale depth datasets, provide superior features for floorplan depth estimation without requiring additional depth training from scratch. From the state-of-the-art models  [5, 16, 55, 20, 52, 53, 6], we select the recent Depth Anything V2 model [53] due to its state-of-the-art performance in both relative and metric depth estimation and low inference times compared to similarly accurate models [6]. Specifically, we use the encoder fine-tuned for indoor environments. We extract features from its last layer and apply bilinear interpolation to match the spatial dimensions the subsequent masked attention mechanism requires.
Let us note that the proposed pipeline is agnostic to the monodepth network, which could easily be replaced as better methods become available. We provide an ablation study on selected encoders in Sec. 4.2.
@@
3.4. Masked Attention. We implement a masked attention mechanism inspired by [10] to predict floorplan depth, using the interpolated features and gravity alignment mask as inputs. Our model predicts two 1D vectors: $\mathbf{\hat{d}}_{t}$, representing floorplan depth estimates, and $\mathbf{\hat{b}}_{t}$, denoting the uncertainty associated with each estimate. The uncertainty quantification allows the model to account for varying confidence levels across regions in the image, particularly in challenging areas with ambiguous visual cues or distant objects.
@@
3.5. Uncertainty-Aware Matching. Using the predicted depth $\mathbf{\hat{d}}_{t}$ and uncertainty $\mathbf{\hat{b}}_{t}$, we compute the observation likelihood over the entire floorplan. Let us use the predicted floorplan depth and uncertainty as our observation $\mathbf{o}_{t}$ and define the following observation model:
$\small p(\mathbf{o}_{t}\mid\mathbf{s}_{t})=\prod_{j=1}^{R}\frac{1}{2\cdot\tilde{b}_{t,j}}\cdot\exp\left(-\frac{|\tilde{d}_{t,j}-d_{j}(\mathbf{s}_{t})|}{\tilde{b}_{j}}\right),$ (1)
where $\tilde{d}_{t,j}$ and $\tilde{b}_{t,j}$ are the predicted depth and uncertainty interpolated from $\mathbf{\hat{d}}_{t}$ and $\mathbf{\hat{b}}_{t}$ at ray angle $\alpha_{j}$, and $R$ is the number of rays. The corresponding floorplan depth $d_{j}(\mathbf{s}_{t})$ is computed from the floorplan ray length as follows:
$\small d_{j}(\mathbf{s}_{t})=r_{j}(\mathbf{s}_{t})\cdot\cos(\alpha_{j}),$ (2)

### v2

@@ Abstract
We propose UnLoc, an efficient data-driven solution for sequential camera localization within floorplans. Floorplan data is readily available, long-term persistent, and robust to changes in visual appearance. We address key limitations of recent methods, such as the lack of uncertainty modeling in depth predictions and the necessity for custom depth networks trained for each environment. We introduce a novel probabilistic model that incorporates uncertainty estimation, modeling depth predictions as explicit probability distributions. By leveraging off-the-shelf pre-trained monocular depth models, we eliminate the need to rely on per-environment-trained depth networks, enhancing generalization to unseen spaces. We evaluate UnLoc on large-scale synthetic and real-world datasets, demonstrating significant improvements over existing methods in terms of accuracy and robustness. Notably, we achieve $2.7$ times higher localization recall on long sequences (100 frames) and $42.2$ times higher on short ones (15 frames) than the state of the art on the challenging LaMAR HGE dataset. Code and materials: https://github.com/matthias-wueest/UnLoc
Figure 1: UnLoc processes an input image sequence to predict the floorplan depth (in meters) and associated uncertainty for each image column. Using these predictions, it generates a probability distribution over potential $\text{SE}(2)$ camera poses and outputs the most likely one (blue arrow). The ground truth pose is also shown (red arrow), overlapped by the predicted pose.
@@ 1 Introduction
Camera localization within indoor environments is a fundamental problem in computer vision, essential for applications in augmented reality and robotics. Accurate localization enables devices to understand their spatial context, facilitating tasks such as navigation, object interaction, and autonomous exploration. Traditional localization methods often rely on pre-built 3D models (Sattler et al., 2011; 2016b; Liu et al., 2017; Sarlin et al., 2019; Panek et al., 2022) or extensive image databases (Schonberger and Frahm, 2016; Arandjelovic et al., 2016; Keetha et al., 2023; Wei et al., 2024), which are storage-intensive and require substantial maintenance, limiting their scalability to new or dynamically changing environments.
@@
Floorplans offer a lightweight and readily available alternative for indoor localization. As 2D representations of spaces, floorplans are easy to obtain (Liu et al., 2018; Yue et al., 2023) and remain unaffected by changes in appearance, such as furniture rearrangements or lighting variations. Recent methods have leveraged floorplans for localization by aligning images with the map (Howard-Jenkins and Prisacariu, 2022; Min et al., 2022; Chen et al., 2024a), enabling devices to localize within new environments as long as a floorplan is available.
@@
Among these, F3Loc (Chen et al., 2024a) has emerged as a recent promising approach for sequential visual floorplan localization, significantly outperforming all previous baselines. It integrates observations over time using a histogram filter, achieving impressive accuracy and computational efficiency on a range of datasets. However, as promising as it is, F3Loc has notable limitations that hinder its practical deployment, which we aim to improve upon in this paper:
@@
Lack of Uncertainty Modeling. F3Loc combines monocular and multi-view depth estimation. However, it assumes that the resulting predictions are all of similar accuracy, and it has no means to represent and predict the uncertainties in the depth predictions. In indoor environments, depth estimation is often unreliable in regions with glass walls, open doorways, or large, featureless walls. When fusing a sequence of predictions, not accounting for uncertainty, inaccurate depth predictions adversely affect the localization process, leading to erroneous pose estimates.
@@
Dataset-Specific Depth. F3Loc relies on a custom depth prediction network trained separately for each dataset or environment. This per-dataset training requirement poses significant challenges for scalability and robustness. Collecting sufficient depth data for retraining the depth network for every new environment is impractical, especially when rapid deployment is desired.
@@
This paper addresses these shortcomings by introducing a novel visual floorplan localization method that incorporates uncertainty estimation into monocular depth prediction and uses it to robustly fuse predictions from a sequence of images. Also, our method enables us to leverage pre-trained monocular depth models. Our contributions are as follows:
@@
Uncertainty-Aware Depth Prediction: We model floorplan depth predictions as explicit probability distributions, assuming a Laplace distribution centered at the predicted depth with scale parameter given by the uncertainty estimate. This formulation allows us to represent the confidence associated with each prediction. This uncertainty-aware approach improves localization by weighting predictions according to their reliability in challenging regions, and it further provides principled weights for the post-processing optimization, yielding higher accuracy.
Leveraging Off-the-Shelf Depth Models: Rather than designing or training custom depth networks for each environment as F3Loc does, we directly employ state-of-the-art monodepth models pre-trained on large-scale datasets (Yang et al., 2024b). Our formulation treats these models as plug-and-play modules, demonstrating that reliable localization can be achieved with any sufficiently strong depth predictor, without requiring environment-specific retraining.
@@
The proposed UnLoc (see Fig. 1) achieves significant improvements in accuracy and robustness across multiple datasets compared to prior methods.
@@ 2 Related work
Visual localization is a fundamental problem in computer vision, addressed through various approaches. Traditional methods include image retrieval techniques (Chum et al., 2007; Jégou et al., 2010; Arandjelovic et al., 2016; Keetha et al., 2023; Wei et al., 2024), which find the most similar images in a database and estimate the pose of the query image based on the retrieved ones. Structure-from-Motion-based approaches (Agarwal et al., 2011; Schonberger and Frahm, 2016; Sattler et al., 2016a; Panek et al., 2022) build a 3D model of the environment and establish 2D-3D correspondences by matching local descriptors, computing camera poses using minimal solvers (Kukelova et al., 2008) and RANSAC (Fischler and Bolles, 1981) or its recent variants (Barath et al., 2020; Barath and Matas, 2021). Scene coordinate regression methods (Brachmann et al., 2017b; a) learn to regress the 3D coordinates of image pixels, while pose regression techniques (Kendall et al., 2015; Kendall and Cipolla, 2017) use networks to predict a 6-DoF camera pose from input images directly. More recently, we saw methods that combine language-based retrieval (Chen et al., 2024b) in combination with scene graph representations (Miao et al., 2024; Zhang et al., 2025). These methods often rely on pre-built 3D models that are storage-intensive and scene-specific, limiting their applicability in unseen environments.
@@
To overcome this, floorplan-based localization methods have emerged, utilizing overhead images or floorplans to estimate the $\text{SE}(2)$ pose of the camera (Workman et al., 2015; Tian et al., 2017). These approaches can localize images in new scenes as long as a floorplan is provided. Floorplan localization is frequently associated with LiDAR sensors (Mostofi et al., 2014; Yin et al., 2019; Zimmerman et al., 2022), which are impractical for widespread mobile device use. Alternative methods reconstruct 3D geometry using depth cameras (Winterhalter et al., 2015) or visual odometry (Mur-Artal et al., 2015). Some approaches extract geometric features like room edges to align with the floorplan (Boniardi et al., 2019; Lin et al., 2019). However, these methods often assume known camera or room height, which is not always feasible.
@@
Recent learning-based methods aim to use only RGB images for floorplan localization. OrienterNet (Sarlin et al., 2023) localizes images in 2D public maps such as OpenStreetMap using neural matching, but focuses on outdoor environments. LaLaLoc (Howard-Jenkins et al., 2021) estimates the position of panoramic images in a floorplan by embedding map and image features into a shared space. LaLaLoc++ (Howard-Jenkins and Prisacariu, 2022) removes the known camera and ceiling height assumption by directly embedding the floorplan. LASER (Min et al., 2022) represents the floorplan as a set of points and uses PointNet (Qi et al., 2017) to embed the visible points for each pose, aligning them with image features in a shared space. PF-Net (Karkus et al., 2018) integrates localization within a differentiable particle filtering framework, using a learned similarity between images and corresponding map patches. F3Loc (Chen et al., 2024a) utilizes metric monocular depth prediction but requires training custom depth networks, which can limit generalizability, and assumes that all predictions are of the same quality.
@@
Sequence-based localization methods enhance robustness by integrating information over time. Bayesian filtering (Dellaert et al., 1999; Chu et al., 2015; Karkus et al., 2018; Boniardi et al., 2019; Mendez et al., 2020) is commonly used to fuse sequential observations, like particle (Dellaert et al., 1999) and histogram filters (Thrun, 2002). PF-Net (Karkus et al., 2018) employs a differentiable particle filter for localization but relies on learned observation models that may not generalize well. F3Loc (Chen et al., 2024a) uses sequential observations and integrates them by a histogram filter.
@@
Depth estimation provides valuable geometric information for localization (Chen et al., 2024a). Recent advances in pre-trained monocular depth (monodepth) models  (Ranftl et al., 2021; Birkl et al., 2023; Guizilini et al., 2023; Yin et al., 2023; Hu et al., 2024; Yang et al., 2024a; b; Bochkovskii et al., 2024), trained on large-scale datasets, enable accurate metric or relative depth predictions out-of-the-box, without per-scene training. In this work, we leverage such off-the-shelf networks for image-based floorplan localization, thereby avoiding custom model training and improving generalization across domains.
Depth uncertainty is crucial for reliable use of depth predictions in downstream tasks. Neural network uncertainty is commonly categorized as aleatoric (data) or epistemic (model) (Kendall and Gal, 2017; Poggi et al., 2020). Aleatoric uncertainty captures inherent observation noise by modeling a distribution over the network output. It is most relevant in regions of the observation space with higher noise and does not decrease with more data. A typical approach trains the network to predict parameters of a parametric distribution via log-likelihood maximization (Nix and Weigend, 1994), usually adding negligible computational overhead (Kendall and Gal, 2017).
Epistemic uncertainty reflects model limitations by placing a distribution over the model parameters. It is most useful with small datasets or safety-critical tasks and decreases as more data becomes available. Estimating epistemic uncertainty generally incurs significant computation cost (Kendall and Gal, 2017). Common strategies include Monte Carlo Dropout (Srivastava et al., 2014), bootstrapped ensembles (Lakshminarayanan et al., 2017), and Bayesian neural networks (MacKay, 1992).
For pixel-wise monocular depth estimation, prior work has explored both types. Kendall and Gal (2017) propose a method to estimate aleatoric and epistemic uncertainty jointly. Poggi et al. (2020) evaluate approaches of both types and propose a self-teaching model for aleatoric uncertainty. Liu et al. (2019) model aleatoric uncertainty by discretizing depth into bins. Roessle et al. (2022) predict aleatoric uncertainty via Gaussian depth distributions with learned variance. In floorplan localization, F3Loc (Chen et al., 2024a) estimates floorplan depth assuming a uniform confidence. In contrast, we model aleatoric uncertainty explicitly to capture ambiguities from glass walls, doorways, and occlusions. Although epistemic uncertainty could help for out-of-distribution scenes, we focus on aleatoric uncertainty for computational efficiency in real-time histogram filtering, leaving epistemic modeling for future work. Our Laplace-based formulation enables uncertainty-aware matching, improving robustness and convergence in challenging indoor environments.
@@ 3 Floorplan Localization with UnLoc
Figure 2: Main method overview. At timestep $t$, UnLoc aligns an image with gravity and processes it through a monodepth encoder. The extracted features, along with a binary mask from the gravity alignment, are used to predict the floorplan depth $\mathbf{\hat{d}}_{t}$ and uncertainty $\mathbf{\hat{b}}_{t}$ via masked attentions. These predictions form equiangular rays, allowing for uncertainty-aware matching with the floorplan’s occupancy map. A histogram filter fuses the observation likelihood with the integrated past belief.
3.1. Method Overview. We estimate the 2D pose $\mathbf{s}_{t}=[s_{x,t},s_{y,t},s_{\phi,t}]$ within a known 2D floorplan, where $s_{x,t}$ and $s_{y,t}$ denote position coordinates and $s_{\phi,t}$ represents orientation. Our approach takes as input a sequence of $t$ RGB images, relative poses between these images, gravity direction, camera intrinsics, and the geometric layout of the floorplan. To ensure out-of-the-box usability of the method, input floorplans are provided solely as occupancy grid data without any semantic labels.
An overview of our main method is shown in Fig. 2. Each image is aligned with gravity and processed by a pre-trained encoder (Yang et al., 2024b) to extract features. These features, along with a binary mask from the gravity alignment, are input to a masked attention mechanism that predicts the floorplan depth, defined as the depth to the nearest occupied area in the floorplan. The predicted depth is then used to construct equiangular rays, which are matched against the floorplan to produce a localization estimate. Finally, we combine this estimate with the previous posterior localization using a histogram filter to compute the updated posterior estimate.
@@
3.2. Gravity Alignment. Real-world applications often involve images captured from hand-held or head-mounted devices, resulting in arbitrary orientations. To address this, we preprocess images to align them with gravity. We achieve alignment by utilizing the camera’s roll ($\psi$) and pitch ($\theta$) angles to define rotation matrices. The rotation for roll is given as $\mathbf{R}_{x}(\psi)$ and for pitch as $\mathbf{R}_{y}(\theta)$. The combined matrix from the gravity-aligned frame to the camera frame is calculated as $\mathbf{R}_{cg}=\mathbf{R}_{y}(\theta)\cdot\mathbf{R}_{x}(\psi)$. The inverse transformation, from the camera frame to the gravity-aligned frame, is given by: $\mathbf{R}_{gc}=\mathbf{R}_{cg}^{\top}$. Using these rotations, we compute a homography $\mathbf{H}$ to warp the image into the gravity-aligned frame $\mathbf{H}=\mathbf{K}\cdot\mathbf{R}_{gc}\cdot\mathbf{K}^{-1}$, where $\mathbf{K}$ is the camera intrinsic matrix. This process also produces a binary mask indicating pixels that are invalid after alignment due to the warping. The gravity-aligned RGB image and the mask are then used for subsequent feature extraction.
@@
Note that the gravity direction and camera intrinsics are usually accessible from sensors on smartphones and head-mounted devices. If this information is not directly available, methods such as GeoCalib (Veicht et al., 2024) can be used to estimate both the intrinsics and gravity direction.
@@
3.3. Feature Extraction. Recent visual floorplan localization methods (Howard-Jenkins et al., 2021; Howard-Jenkins and Prisacariu, 2022; Min et al., 2022; Chen et al., 2024a) rely on encoders pre-trained on ImageNet (Deng et al., 2009) for image classification tasks, such as ResNet-50 (He et al., 2016). However, we posit that encoders trained on tasks more closely related to floorplan depth estimation could yield better performance. To explore this, we utilize encoders pre-trained on dense monocular depth estimation tasks. These encoders, optimized on large-scale depth datasets, provide superior features for floorplan depth estimation without requiring additional depth training from scratch. From the state-of-the-art models  (Birkl et al., 2023; Guizilini et al., 2023; Yin et al., 2023; Hu et al., 2024; Yang et al., 2024a; b; Bochkovskii et al., 2024), we select the recent Depth Anything v2 model (Yang et al., 2024b) due to its state-of-the-art performance in both relative and metric depth estimation and low inference times compared to similarly accurate models (Bochkovskii et al., 2024). Specifically, we use the encoder fine-tuned for indoor environments. We extract features from its last layer and apply bilinear interpolation to match the spatial dimensions the subsequent masked attention mechanism requires. Note that the proposed pipeline is agnostic to the monodepth network, which could easily be replaced as better methods are published.
@@
3.4. Masked Attention. We implement a masked attention mechanism inspired by Chen et al. (2024a) to predict floorplan depth, using the interpolated features and gravity alignment mask as inputs. Our model predicts two 1D vectors: $\mathbf{\hat{d}}_{t}$, representing floorplan depth estimates, and $\mathbf{\hat{b}}_{t}$, denoting the uncertainty associated with each estimate. The uncertainty quantification allows the model to account for varying confidence levels across regions in the image, particularly in challenging areas with ambiguous visual cues or distant objects.
@@
3.5. Uncertainty-Aware Matching. Using the predicted depth $\mathbf{\hat{d}}_{t}$ and uncertainty $\mathbf{\hat{b}}_{t}$, we compute the observation likelihood over the entire floorplan. The predicted uncertainty $\mathbf{\hat{b}}_{t}$ represents aleatoric uncertainty, capturing the inherent observation noise in floorplan depth estimation that arises from scene ambiguities such as glass surfaces, open doorways, and featureless walls. To formulate the observation model, we treat each predicted depth value as drawn from a probability distribution. We model the predicted floorplan depth and its uncertainty as our observation $\mathbf{o}_{t}$ and define the observation likelihood as
$p(\mathbf{o}_{t}\mid\mathbf{s}_{t})=\prod_{j=1}^{R}\frac{1}{2\cdot\tilde{b}_{t,j}}\cdot\exp\left(-\frac{|\tilde{d}_{t,j}-d_{j}(\mathbf{s}_{t})|}{\tilde{b}_{t,j}}\right),$ (1)
where $\tilde{d}_{t,j}$ and $\tilde{b}_{t,j}$ are the predicted depth and uncertainty interpolated from $\mathbf{\hat{d}}_{t}$ and $\mathbf{\hat{b}}_{t}$ at ray angle $\alpha_{j}$, and $R$ is the number of rays. The corresponding floorplan depth $d_{j}(\mathbf{s}_{t})$ is computed from the floorplan ray length as
$d_{j}(\mathbf{s}_{t})=r_{j}(\mathbf{s}_{t})\cdot\cos(\alpha_{j}),$ (2)